OpenStudy (anonymous):

Evaluate f(x) = -2x - 5 for x = 3. A) 1 B) -6 C) 11 D) -11

OpenStudy (anonymous):

@bohotness

OpenStudy (bohotness):

ok first we need to for x

OpenStudy (bohotness):

try put each answer in to see if yuo get the sum you need :P

OpenStudy (anonymous):

f (x) = -2 x - 5 f (3) = -2 (3) - 5 f (3) = -6 - 5 f (3) = -11 This is what I got

OpenStudy (bohotness):

yes thats rigfht :0